One maid can clean the house in 3 hours. Another maid can do the job in 8 hours. How long will it take them to do the job working together?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

One maid can clean the house in 3 hours. Another maid can do the job in 8 hours. How long will it take them to do the job working together?
------------------
1st maid time = 3 hrs/job ; rate = 1/3 job/hr
2nd maid time = 8 hrs/job ; rate = 1/8 job/hr
----
Together time = x hrs/job ; rate = 1/x job/hr
--------------
Equation:
rate + rate = together rate
1/3 + 1/8 = 1/x
------
8x + 3x = 24
11x = 24
x = 24/11 = 2.18 hrs = 2hr + 10min 48sec (time to do the job together)
----------------
Cheers,
Stan H.
--------------------

Answer by ikleyn(52786)

.
A wide variety of similar introductory joint-work problems were solved for you with detailed explanations in these lessons

    - Using Fractions to solve word problems on joint work,
    - Solving more complicated word problems on joint work,
    - Selected joint-work word problems from the archive

Read them and get be trained in solving joint-work problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook under the topic "Rate of work and joint work problems" of the section "Word problems".